# A Deep Learning Method for Simultaneous Denoising and Missing Wedge Reconstruction in Cryogenic Electron Tomography

### v2

@@ Appendix D Details on Experiments
  • Details on our DeepDeWedge: We fitted the model on 418 sub-tomograms until the loss on a hold-out validation set did not improve anymore, which took about 1500 epochs. However, we also obtained reconstructions that were visually similar to the one shown in Figure 3 around the 500 epoch mark.
D.2 Details on Section 2.2.2
  • FBP and sub-tomograms: We followed the steps described in the CryoCARE GitHub repository (https://github.com/juglab/cryoCARE_T2T/tree/master/example) to reconstruct a tomogram from the tilt series with IMOD [KMM96]. We downsampled the tomogram by a factor of 6 using average pooling, which resulted in a voxel size of $14.13$. As IsoNet is optimized for a missing wedge of $60^{\circ}$, we artificially widened the missing wedge from the original $50^{\circ}$ to $60^{\circ}$ by multiplying the reconstruction with a missing wedge filter in the Fourier domain for a fair comparison.
@@
D.3 Details on Section 2.2.3
  • FBP and sub-tomograms:
    We used IMOD to reconstruct the tomograms. First, we binned the pre-aligned tilt series 6 times, which resulted in a pixel size of 20.52 Å. Next, we performed phase flipping based on the provided defocus values for CTF correction, and reconstructed the tomograms via FBP.
    We extracted sub-tomograms with shape $96\times 96\times 96$ for model fitting. Analogous to what we described in Section D.1, we used masks generated with IsoNet to exclude empty areas from sub-tomogram extraction; each sub-tomogram had to contain at least 30% sample according to the mask.
@@
D.4 Details on Section 2.2.4
  • FBP and sub-tomograms: We performed FBP using tomosipo, this time with a Hamming-like filter (see below). We fitted all models for all methods using 150 sub-tomograms of size $96\times 96\times 96$. To refine the FBP reconstructions after model fitting, we applied the final fitted networks to sub-tomograms of the same size, using an overlap of 32 voxels.
@@
Appendix E Further Experiments
If not explicitly stated otherwise, we fitted all models on simulated tilt series from the first 3 tomograms of the Dataset by [Gub+20, ] with SNR $1/4$. We performed only one run for each experiment. All remaining details regarding the model and optimizer are as described in Section 2.2.4 and Appendix D.
E.1 Influence of Sub-Tomogram Size on Performance
Figure 8: Performance metrics for different sub-tomogram sizes used for model fitting.
The size of the sub-tomograms used for model fitting is a hyperparameter of DeepDeWedge, and here we investigate how it affects the performance. To this end, we repeated the experiment on synthetic data described in Section 2.2.4 for increasing the size of sub-tomograms. We scaled the total number of sub-tomograms for each run such that the number of voxels of the sub-tomogram fitting dataset is approximately constant. As for model fitting, we have to rotate the sub-tomograms, the sub-tomograms we extract have to be larger than the ones we actually use for fitting in order to avoid having to use padding. As a result, the maximum sub-tomogram size we could use for model fitting is $128\times 128\times 128$.
E.2 Applying DeepDeWedge to Multiple Tilt Series Simultaneously
As stated in the main body, and as done in some of our experiments in Section 2.2, DeepDeWedge can be applied to multiple tilt series from similar samples simultaneously. Here, we discuss how this impacts performance compared to applying DeepDeWedge to each tilt series separately. We compare the performance of DeepDeWedge on the first 3 tomograms of the synthetic dataset for two scenarios:
@@ E.2 Applying DeepDeWedge to Multiple Tilt Series Simultaneously
Appendix F Uncertainty in Reconstructions of Objects Perpendicular to the Electron Beam
@@
Figure 9 displays an extreme case of a structure that is almost perfectly perpendicular to the electron beam (see blue box). As there is no ground truth, we do not know if the DeepDeWedge prediction of the content inside the blue box is correct or not. Another interesting observation is that the DeepDeWedge reconstruction has relatively high variance across different epochs in this region, which is likely connected to high uncertainty due to missing data.
Figure 9: Slices through reconstructions of Tomogram 3 from EMPIAR-11078. The predictions in the blue box are based on very little data, as the predicted structure in the box is almost perfectly perpendicular to the electron beam. Therefore, it is unclear if the prediction is correct, which is also indicated by the high voxel-wise variance in that area.
Appendix G A Note on Hallucinations
Figure 10: Reconstructions of Tomogram 5 of EMPIAR-10045 obtained with the current version and an old version of DeepDeWedge.
@@ Appendix G A Note on Hallucinations
We believe that the hallucinations shown in Figure 10 result from a mismatch between the model inputs used during model fitting and the inputs used for final reconstruction in the older version of the algorithm. In the older version of DeepDeWedge, the final refinement step, i.e. Step 3 described in Section 2.1, was slightly different than the one presented in Section 2.1: We used the FBP reconstruction of the full, non-splitted tilt series as input to the fitted model. Moreover, we did not progressively fill in the missing wedges of the model inputs during fitting in the previous version. This resulted in said mismatch: During fitting, the model received sub-tomograms from the FBP reconstructions of the splitted tilt series with an additional missing wedge as input. These are noisier and have stronger artifacts than the sub-tomograms of the full FBP reconstruction with only the original missing wedge. In the current version of DeepDeWedge, we try to minimize this mismatch by progressively filling the missing wedge of the model inputs during fitting and refining the FBP reconstructions of the splitted tilt series.
Finally, we note that we observed similar hallucinations as the ones shown in Figure 10 with our re-implementation of Cryo-CARE when applying the fitted model to the FBP reconstruction of the full tilt series instead of applying it to the FBP reconstructions of the even and odd sub-tilt-series and averaging the resulting denoised tomograms as proposed by [Buc+19, ].

### v3

@@ Appendix 4 Details on Main Experiments
  • Details on DeepDeWedge: We fitted the model on 418 sub-tomograms until the loss on a hold-out validation set did not improve anymore, which took about 1500 epochs. However, we also obtained reconstructions that were visually similar to the one shown in Figure 3 around the 500 epoch mark.
4.2 Experiments on Flagella of C. Reinhardtii
The dataset is a single tilt series collected from the flagella of C. reinhardtii. The projections were collected at angles from $-65^{\circ}$ to $+65^{\circ}$ with $2^{\circ}$ increment. Each projection was acquired using dose fractionation with $10$ frames per tilt angle. Each frame has pixel size $2.36$.
  • FBP and sub-tomograms: We followed the steps described in the CryoCARE GitHub repository (https://github.com/juglab/cryoCARE_T2T/tree/master/example) to obtain a CTF-corrected tomogram from the tilt series with IMOD [36]. We downsampled the tomogram by a factor of 6 using average pooling, which resulted in a voxel size of $14.13$. As IsoNet is optimized for a missing wedge of $60^{\circ}$, we artificially widened the missing wedge from the original $50^{\circ}$ to $60^{\circ}$ by multiplying the reconstruction with a missing wedge filter in the Fourier domain for a fair comparison.
@@
4.3 Experiments on the Ciliary Transit Zone of C. Reinhardtii
For our comparison, we chose Tomograms 2, 3, and 8, since their tilt series were approximately collected at the angular range from $-60^{\circ}$ to $+60^{\circ}$ (increment: $2^{\circ}$), for which IsoNet was optimized. We fitted DeepDeWedge and the CryoCARE denoiser for CryoCARE + IsoNet using the frame-based split of the tilt-series data.
  • FBP and sub-tomograms: We used IMOD to reconstruct the tomograms from the raw movie frames. First, we binned the pre-aligned tilt series 6 times, which resulted in a pixel size of 20.52 Å. Next, we performed phase flipping based on the provided defocus values for CTF correction and reconstructed the tomograms via FBP.
    We extracted sub-tomograms with shape $96\times 96\times 96$ for model fitting. Again, we used masks generated with IsoNet to exclude empty areas from sub-tomogram extraction; each sub-tomogram had to contain at least 30% sample according to the mask.
@@
4.4 Experiments on Synthetic Data
  • FBP and sub-tomograms: We performed FBP using tomosipo [31], this time with a Hamming-like filter (see below). We fitted all models for all methods using 150 sub-tomograms of size $96\times 96\times 96$. To refine the FBP reconstructions after model fitting, we applied the final fitted networks to sub-tomograms of the same size, using an overlap of 32 voxels.
@@ 4.4 Experiments on Synthetic Data
Appendix 5 IsoNet Reconstructions with Built-In Denoiser
Supplementary Figure 8: Reconstructions of the C. reinhardtii flagella. IsoNet was fitted with its built-in Noisier2Noise-like denoisier rather than pre-denoising the FBP reconstruction with CryoCARE, resulting in a more noisy reconstruction.
Appendix 6 Further Experiments
If not explicitly stated otherwise, we fitted all models on simulated tilt series from the first 3 tomograms of the SHREC 2021 Dataset with SNR $1/4$. We performed only one run for each experiment. All remaining details regarding the model and optimizer are as in the Methods Section of the main document.
6.1 Applying DeepDeWedge to Multiple Tilt Series Simultaneously
As stated in the main body, and as done in some of our experiments, DeepDeWedge can be applied to multiple tilt series from similar samples simultaneously. Here, we discuss how this impacts performance compared to applying DeepDeWedge to each tilt series separately. We compare the performance of DeepDeWedge on the first 3 tomograms of the SHREC 2021 dataset for two scenarios:
@@ 6.1 Applying DeepDeWedge to Multiple Tilt Series Simultaneously
6.2 Influence of Sub-Tomogram Size on Performance
Supplementary Figure 9: Performance metrics for different sub-tomogram sizes used for model fitting.
The size of the sub-tomograms used for model fitting is a hyperparameter of DeepDeWedge, and here we investigate how it affects the performance. To this end, we repeated the experiment on synthetic data described for increasing the size of sub-tomograms. We scaled the total number of sub-tomograms for each run such that the number of voxels of the sub-tomogram fitting dataset is approximately constant. As for model fitting, we have to rotate the sub-tomograms, the sub-tomograms we extract have to be larger than the ones we actually use for fitting in order to avoid having to use padding. As a result, the maximum sub-tomogram size we could use for model fitting is $128\times 128\times 128$.
Appendix 7 Reconstructing Objects Perpendicular to the Electron Beam
Supplementary Figure 10: Slices through reconstructions of Tomogram 3 from EMPIAR-11078.
@@ Appendix 7 Reconstructing Objects Perpendicular to the Electron Beam
Figure 10 displays an extreme case of a structure that is almost perfectly perpendicular to the electron beam. As the prediction in the blue box is based on very little data, and as there is no ground truth, it is unclear whether the predicted structure is correct. This is also indicated by the high voxel-wise variance over the course of different epochs in that area.
Appendix 8 A Note on Hallucinations
Supplementary Figure 11: Reconstructions of Tomogram 5 of EMPIAR-10045 obtained with the current version and an old version of DeepDeWedge.
@@ Appendix 8 A Note on Hallucinations
We believe that the hallucinations shown in Figure 11 result from a mismatch between the model inputs used during model fitting and the inputs used for final reconstruction in the older version of the algorithm. In the older version of DeepDeWedge, the final refinement step, i.e. Step 3, was slightly different than the one presented in the Results section of the main paper: We used the FBP reconstruction of the full, non-splitted tilt series as input to the fitted model. Moreover, we did not progressively fill in the missing wedges of the model inputs during fitting in the previous version. This resulted in said mismatch: During fitting, the model received sub-tomograms from the FBP reconstructions of the splitted tilt series with an additional missing wedge as input. These are noisier and have stronger artifacts than the sub-tomograms of the full FBP reconstruction with only the original missing wedge. In the current version of DeepDeWedge, we try to minimize this mismatch by progressively filling the missing wedge of the model inputs during fitting and refining the FBP reconstructions of the splitted tilt series.
Finally, we note that we observed similar hallucinations as the ones shown in Figure 11 with our re-implementation of Cryo-CARE when applying the fitted model to the FBP reconstruction of the full tilt series.